Tohban Report 2014-02-19

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 12 Feb 2014
End Date: 19 Feb 2014
Tohban: Hazel Bain
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity was high at the start of the week, with 11 M class flares produced by AR 11974. During the second half of the week solar activity has been low to moderate with the background solar flux sitting around GOES C1 class level. AR 11982 is currently the most active region, producing a number of C class flares over the last 48 hours. There are currently 8 active regions on the disk, 4 of which are approaching the western limb and will rotate off the disk in the next few days, including AR 11974 which produced the M class flares at the start of the week. STEREO B shows several ARs that will rotate onto the disk during the upcoming week.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    57    11     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    39     7     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0    17     5     0

There were RHESSI flares/GOES flares 445 / 68 over the time range 12-Feb-14 19-Feb-14

Memory Management

Front segment data was turned off during eclipse. The change was made at the start of the week on DOY 043 around 20:38UT.

The SSR is now emptying at the end of the daily pass set.

Spacecraft Status

Nominal. The cold plate 1 and 2 are at about 116.7 K and 115.1 K respectively. Cryocooler performance normal.

Data Gaps

N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2014-02-13T18:50:00.000 -- 2014-02-13T19:05:00.000 900.00000 GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2014-02-13T18:50:00.000 -- 2014-02-13T19:00:00.000 600.00000 NO GAPS IN APP_ID = 102 (VC3-MONITOR RATES)

Detector issues

Detector 2 rear resets continue to be seen this week.

Detector 6 and 9 rear fast LLD appears unchanged from last week, remaining similar to that of the other detectors.

Other notes

Updated decimation tables were uploaded.

Spacecraft Management

No operations this week

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Disabled (4 minutes before and after eclipse)
Night time data (rears) Disabled, during entire eclipse
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? None
Personal tools